OpenStudy (anonymous):

X^2+4/x differentiate, confused on how to arrange this

OpenStudy (anonymous):

you have to find the derivative, you use the power rule i just am confused on what goes where,

OpenStudy (phi):

x^2 + 4*x^-1

OpenStudy (phi):

the rule is \[ \frac{d x^n}{dx}= n x^{n-1} \]

OpenStudy (anonymous):

perfect thank you, so you just basically apply the x in the denominator to the outside 4 in this one. I got ya thanks again

OpenStudy (phi):

you should get \[2 x -4x^{-2} \] or \[2 x - \frac{4}{x^2} \]